1. If f'(x) = √2x2 - 1 and y = ƒ(x), then dy/dx at x = 1 is ??

Grade:Upto college level

2 Answers

Mir Noaman Ali IIT-Roorkee
13 years ago

dy/dx is given just put x=1 in that we get f'(1) = root 2 -1

Sudheesh Singanamalla
13 years ago

it is given that f'(x) = √2x2-1

dy/dx = f'(x)

so f'(1) = √2(1)2-1 = √2-1 = 1

 

Final answer is '1'

You can even check this result by applying first principle of differenciation

 

Please approve !